Problem 6.37

In recent years, astronomers have found planets orbiting nearby stars that are quite different from planets in our solar system. Kepler-12b, has a diameter that is 1.7 times that of Jupiter (RJupiter = 6.99
Answer #1
Concepts and reason

The concepts used to solve this problem are acceleration due to gravity and solar system.

Calculate the acceleration due to gravity on the size of Kepler-12b planet.

Fundamentals

The expression for the acceleration due to gravity of the planet is,

g=Gmr2g = \frac{{Gm}}{{{r^2}}}

Here, gg is the acceleration due to gravity of the planet, G is the gravitational constant, m is the mass of the planet, and r is the radius of the planet.

The acceleration due to gravity of the planet is,

g=Gmr2g = \frac{{Gm}}{{{r^2}}}

Rewrite the expression in terms of acceleration due to gravity of Kepler-12b planet.

g=G(0.43mJ)(1.7rJ)2g' = \frac{{G\left( {0.43{m_J}} \right)}}{{{{\left( {1.7{r_J}} \right)}^2}}}

Here, gg' is the acceleration due to gravity of Kepler-12b planet, G is the gravitational constant, mJ{m_J} is the mass of the Jupiter planet, and rJ{r_J} is the means radius of the Jupiter planet.

The expression for the acceleration due to gravity of Kepler-12b planet is,

g=G(0.43mJ)(1.7rJ)2g' = \frac{{G\left( {0.43{m_J}} \right)}}{{{{\left( {1.7{r_J}} \right)}^2}}}

Substitute 6.67×1011m3kg1s26.67 \times {10^{ - 11}}\,{{\rm{m}}^3} \cdot {\rm{k}}{{\rm{g}}^{ - 1}} \cdot {{\rm{s}}^{ - 2}} for GG , 1.90×1027kg1.90 \times {10^{27}}\,{\rm{kg}} for mJ{m_J} , and 6.99×107m6.99 \times {10^7}\,{\rm{m}} for rJ{r_J} to find gg' .

g=(6.67×1011m3kg1s2)(0.43)(1.90×1027kg)[(1.7)(6.99×107m)]2=3.86m/s2\begin{array}{c}\\g' = \frac{{\left( {6.67 \times {{10}^{ - 11}}\,{{\rm{m}}^3} \cdot {\rm{k}}{{\rm{g}}^{ - 1}} \cdot {{\rm{s}}^{ - 2}}} \right)\left( {0.43} \right)\left( {1.90 \times {{10}^{27}}\,{\rm{kg}}} \right)}}{{{{\left[ {\left( {1.7} \right)\left( {6.99 \times {{10}^7}\,{\rm{m}}} \right)} \right]}^2}}}\\\\ = 3.86\,{\rm{m/}}{{\rm{s}}^2}\\\end{array}

Therefore, the acceleration due to gravity of Kepler-12b planet is 3.86m/s23.86\,{\rm{m/}}{{\rm{s}}^2} .

Ans:

The acceleration due to gravity of Kepler-12b planet is 3.86m/s23.86\,{\rm{m/}}{{\rm{s}}^2} .